How do i do this problem? 7h - 11=5h - 7
Answer by checkley79(3341)

7h - 11=5h - 7
7h-5h=-7+11
2h=4
h=4/2
h=2 ans.
Proof:
7*2-11=5*2-7
14-11=10-7
3=3